# Get the Agents Drunk: Memory Perturbations in Autonomous Agent-based Recommender Systems

### v1

@@ 2.1. Victim LLM-powered Agentic RSs
where $\mathcal{X}_{u,v}=\mathcal{P}\oplus m_{u}\oplus\mathcal{M}_{v}^{u}\oplus\mathcal{C}_{v}^{u}$ and $\mathcal{R}_{u}$ is the recommendation results for $u$. $\oplus$ denotes the textual string concatenation. $\mathcal{P}$ stands for prompt templates of recommendations. $m_{u}$ indicates user features and $\mathcal{M}_{v}^{u}=\{m_{v}:v\in\mathcal{V}_{u}\}$ represents the feature set of $u$’s interacted items. $\mathcal{C}_{v}^{u}$ denotes the feature set of candidate items that are not interacted by $u$, where the amount often depends on the type of task (Xu et al., 2021). $|\mathcal{C}_{v}^{u}|>1$ if the task is from implicit feedback such as top-$K$ recommendations, and $|\mathcal{C}_{v}^{u}|=1$ if it is formulated as an explicit feedback task such as rating predictions. In addition to basic textual metadata (e.g., item titles), the features in the vanilla LLM-based RSs include user/item ID information, while the features in the generative agent-based RSs involve the memories of user/item agents. Note that all the elements of $\mathcal{X}_{u,v}$ are optional, which relies on its design.
2.2. Threat Model
2.2.1. Attacker’s Knowledge:
As the works (Ning et al., 2024; Yang et al., 2024; Lin et al., 2022; Zhang et al., 2024d; Zeng et al., 2023) point out, the inherent details (e.g., the architectures, gradients and parameters) of the victim RSs often cannot be obtained in practice. As such, our attack is performed under the black-box settings. In addition, the attacker is only able to access publicly available recommendation data (e.g., item titles and user reviews that can be collected directly by browsing or writing crawlers), due to the openness nature of RSs (Si and Li, 2020; Yang et al., 2023).
2.2.2. Attacker’s Capability:
Unlike another state-of-the-art LLM-powered RS attack (Ning et al., 2024) with an unrealistic assumption of $\mathcal{P}$ and $\mathcal{V}_{u}$ can be easily perturbed (that will be detailed in Section 5), we consider that the attackers have limited perturbation capability in practice. Specifically, they can only interfere with the representation of the target item, which is feasible and can be easily accomplished via a dedicated API, as the existing practical attacks do (Zhang et al., 2024d; Yang et al., 2024; Liu and Larson, 2021).
2.2.3. Attacker’s Objectives:
As mentioned in Section 1, two objectives: transferability and imperceptibility are considered for the practical attack. Given that push attacks are generally used by unscrupulous parties on RSs for profits in practice (Si and Li, 2020) with the aim of promoting target items to as many normal users as possible (Zhang et al., 2021), we thus focus on such attacks in this paper, as normally do in recent works (Wang et al., 2024a; Zhang et al., 2024d). Our first goal is to maximize the hit ratio of the target item $t$ on a surrogate RS (denoted as $\text{Agent}_{\Phi}$ with parameters $\Phi$) that is different from the victim black-box RS, ensuring transferability. Our second goal is to make the descriptions of $t$ as stealthy as possible, achieving imperceptibility. As such, the surrogate RS is used to optimize attacks to avoid frequently querying the victim models to induce suspicion.
Figure 2. Practical Attack Framework DrunkAgent Overview
3. DrunkAgent
In this section, we present a practical attack framework, DrunkAgent, for gaining insights into the robustness of autonomous agent-based RSs with powerful LLMs and related defense models in practice. DrunkAgent is composed of three attack modules. Fig.  2 gives an overview of DrunkAgent, where each component will be described in the following.
3.1. Surrogate Module
We leverage a surrogate RS $\Phi$ to measure how effective the attack is under the black-box settings, as inspired by the existing works (Lin et al., 2022; Tang et al., 2020; Yang et al., 2023, 2024). It is used to optimize the generation module and the strategy module, making the attack transferable and imperceptible.
Since our attack targets SOTA LLM-powered RSs, we adopt powerful LLMs as recommendation agents. Given the transparency, flexibility and scalability of the open-source LLM models, they can be considered for building this module. Moreover, due to our attack focus on adversarial text description generation tasks, we consider an autoregressive language model trained with a causal language modeling objective (Radford et al., 2019; Brown et al., 2020). There are some notable models, such as LLaMA (Touvron et al., 2023a), LLaMA 2 (Touvron et al., 2023b) and LLaMA 3 (Dubey et al., 2024), BLOOM 176B (Le Scao et al., 2023) and Falcon 180B (Almazrouei et al., 2023). Currently, Meta’s LLaMA family has become the most influential open LLM and has been widely adopted in the recent recommendation community (Zhao et al., 2023; Bao et al., 2023). To ensure the effectiveness of attacks, we apply the latest generative model: LLaMA 3, so that valuable feedback can be provided even with limited available resources. Specifically, LLaMA 3 uses an optimized transformer architecture (Vaswani et al., 2017). Its instruction-tuned version uses supervised fine-tuning and reinforcement learning from human feedback (RLHF) to align with human preferences for helpfulness and safety (AI, 2024), as shown in Fig.  2.
In addition to conducting a strict selection of recommendation agents, we develop the two modules for effective target item’s adversarial description generation.
3.2. Generation Module
In this section, we aim to generate adversarial description triggers on the agent-powered RSs, which can promote the target items to as many normal users as possible. The triggers should have three critical properties: 1) Universality on any users and items; 2) Transferability among different black-box victim models; and 3) Imperceptibility to not easier to be discovered. To this end, we design sub-modules, as shown in Fig.  2 and described as follows.
3.2.1. Adversarial Prompting Engineering
Recently, many of the mainstream LLMs that have been widely used in the agentic recommendation community (Huang et al., 2024) (e.g., GPT-3.5-turbo, ChatGPT) are closed source, where the models’ underlying codes, weights, or other critical components are not released to the public and the models can only be invoked through their APIs. As a result, the agent-powered RSs generally adopt in-context learning (ICL) (Dong et al., 2022) to sufficiently take advantage of the generalization capability of LLMs on unknown tasks. To this end, this module carefully crafts prompts $\widetilde{\mathcal{X}_{t}}$ within the ICL framework to improve the quality of adversarial trigger generation.
Basic Prompt Template. We meticulously prepare prompt templates $\widetilde{\mathcal{P}}$ to provide context, goals, instructions and constraints so that the model knows ‘what needs to be done’ and ‘how to respond to’. The main components are described in the following.
  1. (1)
    Style Definition $\widetilde{p_{s}}$. There are two common styles of prompts: task-oriented and role-oriented (Zhu et al., 2023a). To make recommendations more authentic and reasonable, we introduce role information into the context, which allows the system to respond with domain-specific knowledge, such as ‘You are a CD enthusiast.’, as shown in Fig.  2.
  2. (2)
    Task Goal $\widetilde{p_{g}}$. Since a recommender system is essentially a ranker (Zhang et al., 2019), we set sorting goals to indicate recommendation tasks, such as ‘Now, you are considering to sort ten candidate CDs in order of how well they align with your preferences.’ in the given example of the figure.
  3. (3)
    Recommendation Instruction $\widetilde{p_{i}}$. We introduce highly detailed instructions to guide the model to perform accurate recommendations. For example, ‘Please sort CDs by considering the correlation between your preferen -ces/dislikes and the features of the CDs. The higher the CD rank, the more you like it.’ in the shown example.
  4. (4)
@@ 3.2.1. Adversarial Prompting Engineering
Hence, the adversarial prompt template can be formulated as:
(2) $\widetilde{\mathcal{P}}=\widetilde{p_{s}}\oplus\widetilde{p_{g}}\oplus\widetilde{p_{i}}\oplus\widetilde{p_{f}}.$
To increase the attack effectiveness, we introduce position bias (Zhang et al., 2024b) into $\widetilde{\mathcal{P}}$, where the other candidates items are placed before the target item, as shown in the example of Fig.  2.
General User Memories. It is difficult to obtain the optimized memories of user agents (i.e., language-based parametric embeddings of the agentic RSs (Zhang et al., 2024b)), due to the practical attack is launched under the black-box settings. Moreover, the adversarial description triggers are designed to be effective for general users, rather than specific user groups (e.g., users who are only obsessed with Progressive Metal music), so that the target items can be promoted to as many users as possible. Therefore, to enhance the universality of the triggers on users, we use primitive and general user descriptions as user memories $\widetilde{m_{u}}$, such as ‘I enjoy listening to CDs very much.’ in the given example, as the default user memories in the existing works (Zhang et al., 2024b)(Huang et al., 2024).
Popular Item Memories. Given that popular items usually have a high probability of appearing at the top of the users’ recommendation lists, Bandwagon attack (Si and Li, 2020) introduces popular items into fake user profiles to poison traditional RSs for increasing the popularity of the target item and Expert attack on visually-aware RSs (Liu and Larson, 2021) creates adversarial images by decreasing the representation distance to popular items for improving the exposure rate of the target item. Inspired by this, we select popular items as candidate items for $\widetilde{\mathcal{X}_{t}}$ to introduce popularity bias (Zhang et al., 2024b) to further enhance the effectiveness of adversarial description triggers. The basic idea is that the target items should be promoted to more normal users if they can be ranked ahead of the popular items. Let $\mathcal{U}_{v}=\{u\in\mathcal{U}:y_{u,v}\neq 0\}$ indicates the set of users that have interacted with $v$ and $\mathcal{M}_{p}$ denotes a sequence of the features of popular items,
(3) $\mathcal{M}_{p}=\{m_{v}\},$
where $m_{v}$ is the features of the item $v$ including titles and descriptions. $\{v\}_{v\in\mathcal{V}}$ is a sequence of popular item IDs, which is corresponding to the subscript of the item popularity sequence in descending order $\{|\mathcal{U}_{v}|\}$. The item popularity is item frequency here, due to this work focuses on an implicit task.
The descriptions of candidate items will greatly influence the feedback of the surrogate model. As with optimized user agent memories, the adversary is hard to access to internal memories of item agents in practice and is restricted to leverage publicly available data (e.g., item titles and descriptions). However, the raw descriptions (Ni et al., 2019) are often too long that may exceed the context window limitations of the surrogate model and always contain too much redundant information that may increase the difficulty of subsequent optimizations. Similar to the use of user memories, we can consider the default item memories for popular items. Nevertheless, the default ones generally only include identity information such as titles and categories (Zhang et al., 2024b), as the example “The CD is called ‘Brainwashed’. The category of this CD is: ‘Classic Rock; Album-Oriented Rock (AOR)’.” shows, which are overly simple and information-less, leading to the target items easier to be ranked at the top and the sub-optimal results will be produced. To tackle these issues, we propose to adopt a LLM with frozen parameters $\Psi$ to extract valuable item features from the raw descriptions and supplement the features based on its internal rich knowledge base to make them more informative to facilitate the optimizations. Let $\mathcal{C}_{p}$ represent the modified feature sequence of popular items,
(4) $\mathcal{C}_{p}=f_{\text{LLM}_{\Psi}}(\mathcal{X}_{p}\oplus\mathcal{M}_{p}),$
where $\mathcal{X}_{p}$ denotes the prompts for feature acquisition and OpenAI’s GPT-4 (Achiam et al., 2023) is adopted to implement the LLM, because it is a state-of-the-art model with powerful language understanding, reasoning and generation capabilities. An example of producing a popular item memory is given in Fig.  2. To sum up, the adversarial prompt can be formulated as
(5) $\widetilde{\mathcal{X}_{t}}=\widetilde{\mathcal{P}}\oplus\widetilde{m_{u}}\oplus\mathcal{C}_{p}\oplus\widetilde{m_{t}},$
where $\widetilde{m_{t}}$ is the features of the target item. In addition to the fixed title, the adversarial description trigger (i.e., the memory of the item agent) will be optimized in the next sub-section.
Note that we give lightweight and efficient adversarial prompts for the surrogate RS that are different from the real victim RSs’. To ensure high-quality recommendations, more powerful and context-rich (e.g., retrieval-augmented) prompts will be used in the victims, which will be described in the evaluation section.
3.2.2. Greedy Search Algorithm
We propose a greedy search algorithm to optimize the adversarial description triggers, as inspired by the successes of the current jailbreak attacks on LLMs (Zou et al., 2023; Liu et al., 2024b). However, there are some essential differences (e.g., attack goals and victim models) between jailbreak attacks and our attacks. Jailbreak attacks are designed to undermine the LLMs’ human-aligned values or other constraints imposed by model developers, compelling them to respond to malicious questions posed by adversaries with correct answers rather than rejection, whereas our attacks aim to promote the target items on the RSs powered by autonomous agents.
Adversarial Trigger Initialization. Given that LLMs should understand LLM-powered agents better, we leverage a LLM to generate some description candidates. To narrow search space by a large margin to increase the efficiency and effectiveness of the algorithm, the basic attack goals are incorporated into the prompts. In addition, the diversity and the generalizability are introduced to the prompts to help search for the global optimal solution and improve their universality on items, respectively. Formally,
(6) ~$\widetilde{\mathcal{M}}_{c}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{c}}\oplus\widetilde{m_{t}}),$
where ~$\widetilde{\mathcal{M}}_{c}=\{\widetilde{m}_{t_{i}}\}$ represents a sequence of trigger candidates obtained from the LLM and the state-of-the-art GPT-4 is used to implement the LLM. $\widetilde{\mathcal{X}_{c}}$ denotes a sequence of prompt templates. An example of the initialization is given in Fig.  2.
Adversarial Trigger Optimization. The algorithm takes $\mathcal{E}$ epochs to obtain the optimal trigger, where the output candidates of each epoch will be the inputs of the next epoch for continuous optimization until the attack objective is achieved or the termination condition is reached (e.g., the maximum number of epochs is exhausted). To avoid frequently querying the victim models to induce suspicion and mimic the black-box settings in the real world, the optimization is built on the surrogate module, as shown in Fig.  2.
For each epoch $k$, the algorithm first calculate the performance score for each input candidate $\widetilde{m}_{t_{i}}^{(k-1)}$, where $n$ well-performing trigger candidates will be maintained directly for the next optimization to ensure the stability. The higher the score, the more likely the candidate is to be kept. The score is calculated by the negative of the loss, i.e., $s_{i}^{(k-1)}=-\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}$. The smaller the loss, the more likely the target item is to be ranked high and the more effective the candidate is. We use the auto-regressive language generation loss (i.e., negative log-likelihood) to evaluate the discrepancy between the predictions and the target output $\widetilde{\mathcal{R}_{t}}$,
(7) $\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}=-\log\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)})),$
@@ 3.2.2. Greedy Search Algorithm
(8) $\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)}))=\prod_{l=1}^{L}\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{r_{t_{l}}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)}\oplus\widetilde{r_{t_{<l}}})),$
Then, the algorithm samples a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates by conducting probability-based random selection, where the candidates with higher scores will have a greater probability of being selected into this subset. The selection probability of each candidate is calculated by a softmax function:
(9) $\text{softmax}(\widetilde{m}_{t_{i}}^{(k-1)})=\frac{e^{s_{i}^{(k-1)}}}{\sum_{j=1}^{|\widetilde{\mathcal{M}}_{c}|}e^{s_{j}^{(k-1)}}}.$
For the obtained subset, the algorithm splits the candidates at random positions into a list of text slices and performs the random exchange of the slices between pairwise texts. For example, for the pairwise $<\widetilde{m}_{t_{i}}^{(k-1)},\widetilde{m}_{t_{j}}^{(k-1)}>$, $\widetilde{m}_{t_{i}}^{(k-1)}=[\widetilde{m}_{t_{i_{1}}}^{(k-1)},\widetilde{m}_{t_{i_{2}}}^{(k-1)},\widetilde{m}_{t_{i_{3}}}^{(k-1)}]$ and $\widetilde{m}_{t_{j}}^{(k-1)}=[\widetilde{m}_{t_{j_{1}}}^{(k-1)},\widetilde{m}_{t_{j_{2}}}^{(k-1)},\widetilde{m}_{t_{j_{3}}}^{(k-1)},\widetilde{m}_{t_{j_{4}}}^{(k-1)}]$, the new candidates could be $\widetilde{m}_{t_{i}}^{(k-1)^{\prime}}=[\widetilde{m}_{t_{i_{1}}}^{(k-1)},\widetilde{m}_{t_{j_{2}}}^{(k-1)},\widetilde{m}_{t_{i_{3}}}^{(k-1)},\widetilde{m}_{t_{j_{4}}}^{(k-1)}]$ and $\widetilde{m}_{t_{j}}^{(k-1)^{\prime}}=[\widetilde{m}_{t_{j_{1}}}^{(k-1)},\widetilde{m}_{t_{i_{2}}}^{(k-1)},\widetilde{m}_{t_{j_{3}}}^{(k-1)}]$. As a result, the new combinations of different words, phrases, and sentences are gained, which increases the feature diversity of candidates through extensive exploration, contributing to search for the optimal trigger.

### v2

@@ 2.1 Victim LLM-powered Agentic RSs
where $\mathcal{X}_{u,v}=\mathcal{P}\oplus m_{u}\oplus\mathcal{M}_{v}^{u}\oplus\mathcal{C}_{v}^{u}$ and $\mathcal{R}_{u}$ is the recommendation results for $u$. $\oplus$ denotes the integration of textual strings, encompassing both concatenation and interpolation of substrings. $\mathcal{P}$ stands for prompt templates of recommendations. $m_{u}$ indicates user features and $\mathcal{M}_{v}^{u}=\{m_{v}:v\in\mathcal{V}_{u}\}$ represents the feature set of $u$’s interacted items. $\mathcal{C}_{v}^{u}$ denotes the feature set of candidate items that are not interacted by $u$, where the amount often depends on the type of task (xu2021understanding, ). $|\mathcal{C}_{v}^{u}|>1$ if the task is from implicit feedback such as top-$\mathcal{K}$ recommendations, and $|\mathcal{C}_{v}^{u}|=1$ if it is formulated as an explicit feedback task such as rating predictions. In addition to basic textual metadata (e.g., item titles), the features in the vanilla LLM-based RSs include user/item ID information, while the features in the generative agent-based RSs involve the memories of user/item agents. Note that all the elements of $\mathcal{X}_{u,v}$ are optional, which relies on its design.
2.2 Threat Model
2.2.1 Attacker’s Knowledge:
It is often challenging to obtain the internal knowledge of real-world victim RSs, as noted by prior works (ning2024cheatagent, ; yang2024attacking, ; zhang2024stealthy, ). Specifically, critical details such as the parameters (e.g., agents’ memories, which are language-based embeddings (zhang2024agentcf, )) and architectures (e.g., the underlying LLM backbones) are typically inaccessible to adversaries. Given these limitations, DrunkAgent is primarily designed and evaluated under black-box settings (in Section 4). Moreover, due to the inherently open nature of RSs (si2020shilling, ; 10415763, ), in both cases, the attacker is assumed to only have access to publicly available recommendation data such as item titles and user reviews.
2.2.2 Attacker’s Objectives:
As outlined in Section 1, the core objective of our attack is to maximize the exposure of a target item across the widest possible range of normal users. Since black-box attacks are typically optimized in a localized manner (e.g., via surrogate models (wang2024poisoning, )), achieving this objective requires strong transferability, i.e., the ability of the attack to remain effective across different black-box RSs (ning2024cheatagent, ). Furthermore, to broaden the attack impact and maximize the number of affected users in practice, stealthiness becomes a critical secondary objective, ensuring that the attack remains inconspicuous thus difficult to detect (zhang2024stealthy, ).
2.2.3 Attacker’s Capabilities:
In the black-box scenarios, the attacker has no access to the model’s internal states (e.g., user memories and candidate item memories) and can induce memory corruptions by modifying the description of the target item (i.e., the initial memory of the target item agent). Descriptions are often both lengthy and context-rich, which can easily conceal perturbations, as elaborated in Appendix A. Moreover, the profit-driven merchant may frequently update the descriptions to overwrite the target memory that has been optimized over time or construct multiple descriptions of items indicating the same target item. This is feasible, due to the merchants are required to maintain their dedicated APIs (liu2021adversarial, ; yang2024attacking, ; zhang2024stealthy, ).
Figure 2: DrunkAgent Overview. The generation module produces adversarial textual triggers for promoting target items. The strategy module creates adversarial strategies to ‘get the target agents drunk’ to allow the triggers to achieve maximum impact. The triggers and the strategies are optimized on the surrogate module to improve the transferability and stealthiness of black-box attacks.
3 DrunkAgent
In this section, we present the black-box attack framework, DrunkAgent. Fig.  2 gives an overview of DrunkAgent, which is composed of three modules. Each module is introduced in the following.
3.1 Surrogate Module
This module serves as a surrogate agentic RS (denoted as $f_{\text{Agent}_{\Phi}}(\cdot)$) to evaluate attack effectiveness under limited accessible resources, drawing inspiration from prior works lin2022shilling ; tang2020revisiting ; 10415763 ; yang2024attacking . To avoid frequent queries that could arouse suspicion and to adapt to realistic black-box conditions, the generation module and the strategy module are optimized within this surrogate setting to improve the stealthiness and transferability of DrunkAgent. As discussed in Section 1, agentic RSs are composed of user agents, item agents, and recommendation agents, and all built upon LLM backbones enhanced with memory modules, where the LLMs are used to perform text generation tasks. As such, open-source and auto-regressive LLMs are preferable for constructing surrogate backbones due to their strong generation capabilities and customizability. Given that the most influential Meta’s LLaMA family has been widely adopted in the recent recommendation community (zhao2023recommender, ; bao2023tallrec, ), the latest and powerful LLaMA 3 (dubey2024llama, ) is an ideal choice, which can provide more valuable and instructive feedback for effective optimizations, as illustrated in Fig. 2. The memories for each types of surrogate agents will be detailed in the following sub-sections.
3.2 Generation Module
This module aims to generate effective adversarial description triggers for promoting the target items on the agentic RSs, consisting of several sub-modules, as shown in Fig. 2 and described as follows.
3.2.1 Adversarial Prompting Engineering
This module designs the adversarial prompt for the surrogate recommendation agent to improve the quality of adversarial trigger generation. We craft prompt template, so that the surrogate knows ‘what needs to be done’ and ‘how to respond to’. As discussed in Section 2.2.1, it is challenging to obtain optimized memories of user agents and item agents in black-box attacks. To solve this and ensure the quality, we introduce general user memories and popular item memories.
Basic Prompt Template. The template $\widetilde{\mathcal{P}}$ is built with four components: Style Definition $\widetilde{p_{s}}$, Task Goal $\widetilde{p_{g}}$, Recommendation Instruction $\widetilde{p_{i}}$ and Format Constraint $\widetilde{p_{f}}$, as shown in the example of Fig. 2. The detailed descriptions are provided in Appendix B.1. Formally, $\widetilde{\mathcal{P}}=\widetilde{p_{s}}\oplus\widetilde{p_{g}}\oplus\widetilde{p_{i}}\oplus\widetilde{p_{f}}.$
General User Memories. The triggers aim to be effective for general users, rather than niche user groups (e.g., those exclusively interested in Progressive Metal), to maximize the exposure rate of the target items. Hence, the user memory $\widetilde{m_{u}}$ is set with universal user descriptions, as the example shows in Fig. 2. This is also the initial memory of the user agent (zhang2024agentcf, ; huang2024foundation, ), which is accessible and aligns with the assumptions of target black-box RSs.
Popular Item Memories. Besides the limited access of the optimized memories in practice, the inherent opacity of the victim systems renders the exact set of candidate items used for recommendations unknown. As the works (si2020shilling, ; liu2021adversarial, ; wang2025id, ) pointed out, popular items usually have high probabilities of appearing at the top of users’ recommendation lists. Intuitively, if target items can be ranked ahead of these popular items, they are more likely to be promoted to normal users. Therefore, we adopt popular items as candidate items to enhance the effectiveness of trigger optimizations. These popular items are publicly available, which are not beyond the black-box assumptions. Let $\mathcal{U}_{v}=\{u\in\mathcal{U}:y_{u,v}\neq 0\}$ indicates the set of users that have interacted with $v$ and $\mathcal{M}_{p}$ denotes a sequence of the features of popular items. Thus, $\mathcal{M}_{p}=\{m_{v}:v\in\mathcal{V}_{p}\}$, where $\mathcal{V}_{p}$ is a sequence of popular item IDs. The sequence is the corresponding subscript of the item popularity sequence (i.e., $\{|\mathcal{U}_{v}|\}_{v\in\mathcal{V}}$) in descending order. Raw item descriptions serve as a form of natural resource that can be leveraged to derive memories for popular items. However, such descriptions (ni2019justifying, ) are often too long that may exceed the context window limitations of the surrogate model and always contain too much redundant information that may increase the difficulty of subsequent optimizations. To tackle the issues, we propose to adopt a LLM (denoted as $f_{\text{LLM}_{\Psi}}(\cdot)$) to extract valuable item features from the raw descriptions and supplement the features based on its internal rich knowledge base to make them more informative to facilitate the optimizations. Let $\mathcal{C}_{p}$ represent the modified feature sequence of popular items, $\mathcal{C}_{p}=f_{\text{LLM}_{\Psi}}(\mathcal{X}_{p}\oplus\mathcal{M}_{p})$, where $\mathcal{X}_{p}$ denotes the prompts for feature acquisition. An example of producing a popular item memory is given in Fig. 2. Note that if the raw item descriptions are empty, we adopt other identity information including titles and categories.
Agents always tend to select candidates positioned higher in the display list (zhang2024agentcf, ). To increase the trigger effectiveness, we place the other popular candidates before the target item, as shown in Fig. 2. In the previous sections, $m_{v}$ is commonly used to represent a set of features of item $v$. For better understanding the subsequent methods, for the target item $t$, we use different notations to distinguish between meta features and memories: $m_{t}$ denotes the metadata, while $\widetilde{m_{t}}$ denotes the memory (i.e., the adversarial trigger). As a result, the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ can be formulated as
$\widetilde{\mathcal{X}_{t}}=\widetilde{\mathcal{P}}\oplus\widetilde{m_{u}}\oplus\mathcal{C}_{p}\oplus m_{t}.$ (2)
3.2.2 Greedy Search Algorithm
The algorithm is used to optimize the adversarial description triggers, inspired by the recent successes in jailbreak attacks on general-purpose LLMs (zou2023universal, ; liu2023autodan, ). Despite this inspiration, DrunkAgent fundamentally differs in key aspects such as attack goals and victim models. Additional details are given in Appendix A. To fill the gaps, the algorithm incorporates task-aligned initialization and refines the search space tailored to manipulate the behaviors of the target agents.
Adversarial Trigger Initialization. Given that LLMs should understand LLM-powered agents better, we leverage a LLM to generate some descriptive candidates. To narrow search space by a large margin to increase the efficiency and effectiveness of the algorithm, the basic attack goals are incorporated into the prompts. In addition, the diversity and the generalizability are introduced into the prompts to help search for the global optimal solution, as the example of the initialization in Fig. 2 shows. Let $\widetilde{\mathcal{X}_{c}}$ denote a sequence of prompts,
~$\widetilde{\mathcal{M}}_{c}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{c}}\oplus m_{t}),$ (3)
where ~$\widetilde{\mathcal{M}}_{c}=\{\widetilde{m}_{t_{i}}\}$ represents a sequence of trigger candidates obtained from the LLM.
Adversarial Trigger Optimization. The algorithm takes $\mathcal{E}$ epochs to obtain the optimal trigger, where the output candidates of each epoch will be the inputs of the next epoch for continuous optimization until the attack objective is achieved or the termination condition is reached. For each epoch $k$, there are three stages.
Quality Estimation Stage. The algorithm calculates the performance score for each input candidate $\widetilde{m}_{t_{i}}^{(k-1)}$. $n$ well-performing trigger candidates will be maintained directly for the next optimization to ensure the stability. The higher the score, the more likely the candidate is to be kept. The score is calculated by the negative of the loss, i.e., $s_{i}^{(k-1)}=-\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}$. The smaller the loss, the more likely the target item is to be ranked high and the more effective the candidate is. We use the auto-regressive language generation loss (i.e., negative log-likelihood) to evaluate the discrepancy between the predictions and the target output $\widetilde{\mathcal{R}_{t}}$,
$\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}=-\log\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)})),$ (4)
@@ 3.2.2 Greedy Search Algorithm
Feature Integration Stage. The algorithm samples a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates by conducting probability-based random selection. The candidates with higher scores will have a greater probability of being selected, where the selection probability of each candidate is calculated via a softmax function, i.e., $\text{softmax}(\widetilde{m}_{t_{i}}^{(k-1)})=\frac{e^{s_{i}^{(k-1)}}}{\sum_{j=1}^{|\widetilde{\mathcal{M}}_{c}|}e^{s_{j}^{(k-1)}}}$. For the obtained subset, the algorithm splits the candidates at random positions into a list of text slices and performs the random exchange of the slices between pairwise texts, allowing features from different candidates to be fused, as shown in Fig. 2. Specifically, this fusion creates new combinations of words, phrases, and sentences, enhancing the feature diversity of candidates, and hence contributing to search for an optimal trigger.
@@
The optimal trigger $\widetilde{m_{t}}^{*}$ will be obtained when the candidate performs the best,
$\widetilde{m_{t}}^{*}=\mathop{\arg\max}_{\widetilde{m}_{t_{i}}^{\mathcal{E}}}(\{s_{i}^{\mathcal{E}}\}),$ (5)
where $\{s_{i}^{\mathcal{E}}\}$ is the sequence of final candidates’ scores. An example of $\widetilde{m_{t}}^{*}$ is given in Fig. 2.
3.3 Strategy Module
We introduce the strategy module to impede the target item agent from evolving from the environment, that is, the memory of the target agent cannot be effectively updated during agent-environment interactions. As such, the well-crafted adversarial description trigger $\widetilde{m_{t}}^{*}$ is retained, allowing the promotion of the target item to be maximized.
3.3.1 Perturbation Strategy Definition.
[…]
$\widetilde{q}=\widetilde{q_{f}}\oplus\widetilde{q_{c}}\oplus\widetilde{q_{g}}\oplus\widetilde{q_{n}}\oplus\widetilde{q_{s}},$ (6)
where an example of $\widetilde{q}$ is given in Fig. 2.
3.3.2 Adversarial Strategy Optimization.
The arrangement of the defined strategies needs to be optimized to improve the attack effectiveness. DrunkAgent obtains the optimal strategy $\widetilde{q}^{*}$ by maximizing the expected probability that the agent when influenced by adversarial modifications, performs a malicious action for a given input query,
$\mathbbm{E}_{\widetilde{q}\sim\pi_{\widetilde{q}}}[\ \mathbbm{I}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{X}_{s}}\oplus\widetilde{q}\oplus\widetilde{m_{t}}^{*})=\widetilde{a_{t}})],$ (7)
where $\pi_{\widetilde{q}}$ denotes the distribution of adversarial strategies, $\mathbbm{I}(\cdot)$ is an indicator function and $\widetilde{\mathcal{X}_{s}}$ is an adversarial prompt for simulating memory updates of item agents during the agents’ environmental interactions. $\widetilde{a_{t}}$ is the desired malicious action for the injected strategy $\widetilde{q}$, i.e., the target item agent is unable to fulfill specific roles and is incapable of perceiving, learning and self-evolving from the interaction environment (e.g., fail in language-based signal back-propagation (zhang2024agentcf, )). An example of $\widetilde{q}^{*}$ is given in Fig. 2. To sum up, the adversarial description of the target item is $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$. The overall optimization procedure of DrunkAgent will be provided in Algorithm 1.
The above design further reveals how memory mechanisms in agent-based recommender systems can be influenced under black-box conditions. These insights may support future efforts toward developing memory-aware defenses (e.g., using customized deep neural network detectors (wu2020densely, ; yang2021dualnet, ; yang2021hunter, ; yang2021deep, )) and enhancing the robustness of agent-powered RSs chua2024ai .
4 Experiments
4.1 Experimental Setup
Datasets Selection. To comprehensively evaluate DrunkAgent, we use three datasets that are text-intensive containing real-world data and widely-adopted in the current recommendation studies (huang2024foundation, ; zhang2024agentcf, ). They are CDs & Vinyl, Office Products and Musical Instruments of Amazon Review Data (ni2019justifying, ). Their statistics are given in Appendix C.1.1. We additionally evaluate our method on Yelp dataset [^2] https://www.yelp.com/dataset to demonstrate its generalizability across domains in the appendix. The target items are randomly sampled from the item pool on each dataset. The universality of attacks across different target items is further demonstrated in Appendix C.2.1.
Baseline Attack Methods. Since we are the first work to gain insight into the robustness of the generative agent-powered RSs, there is no any baselines. To comprehensively evaluate the effectiveness of DrunkAgent, we follow the state-of-the-art attack on LLM-based RSs (zhang2024stealthy, ) and compare DrunkAgent with the benign status and six black-box attack methods, where the perturbations range from character-level: DeepwordBug (gao2018black, ) and PuncAttack (formento2023using, ), word-level: TextFooler (jin2020bert, ) and BertAttack (li2020bert, ) to sentence-level: TrivialInsertion and ChatGPTAttack (zhang2024stealthy, ). To maintain a fair comparison, the methods perform perturbations on the descriptions of the target items and are optimized by the surrogate model that is the same as DrunkAgent. More detailed implementations are provided in Appendix C.1.2.
Targeted Recommender Systems. We adopt three different agentic paradigms (zhang2024agentcf, ) as victim RSs to evaluate the effectiveness of DrunkAgent across diverse agent designs and recommendation tasks: AgentCF, a standard CF paradigm; AgentRAG, which augments the agents with retrieval; and AgentSEQ, which captures temporal dynamics in user behaviors for sequential recommendations. Given that the victim RSs are totally different from as well as more powerful and context-rich than the surrogate, a strict black-box setting is maintained for sufficient evaluations. To ensure a fair evaluation, we follow the default implementations, as elaborated in Appendix C.1.3.
The configuration of DrunkAgent is given in Appendix C.1.4. We evaluate its performance using widely-used HR@$\mathcal{K}\uparrow$, NDCG@$\mathcal{K}\uparrow$, and sentence perplexity score $\downarrow$. The specific definitions of the metrics are in Appendix C.1.5. More experiments and demonstrations (e.g., ablation studies and the sensitivity of critical parameters) can be found in the appendix.